OpenStudy (anonymous):

-4+r ___________= -1 -8

OpenStudy (anonymous):

r = 12

OpenStudy (anonymous):

separate it to -4/-8 + r/-8 =-1 1/2 + r/-8 =-1 r/-8=-3/2 r=-3/2*-8 =12